# Frostman random variables, entropy inequalities, and applications

### v4

@@ 1 Introduction
These applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be viewed as consequences of the Falconer distance problem. If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress. Given this line of research, it is reasonable to expect different proofs of Theorems 1.51.7 with other methods. In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example. When $\phi(x,y)=x^{2}+y^{2}$ ($\alpha=\beta=1$) and $G=A\times A$, after posting this paper to arXiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative lower bound for the following estimate
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(A^{\wedge 2}+A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
@@ 2.1 Notations
In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$.
All logarithms we use here have base $2$. Moreover, we denote $\log_{+}x:=\max\{\log x,\,0\}$ for any $x>0$.
@@ 2.2 Entropy and basic properties
In this definition, we use the convention $0\log 0=0$. We also interpret $\mu(B)H(\mu_{B},\mathcal{A})$ as 0 when $\mu(B)=0$, even though $\mu_{B}$ is not defined in this situation. Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. Before stating the first lemma, we introduce the following notation:
@@
Moreover, the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$. It is increasing on $0,e^{-1}]$, and decreasing on $e^{-1},1]$. With this in tow, we record some properties of entropy for partitions, whose proofs are simple exercises.
@@
(vi) For any Borel set $Y$, we have $\mu(Y)H(\mu_{Y},\mathcal{A})\leq H(\mu,\mathcal{A}).$
@@
For the random variable $Z=(X,Y)$, we denote $H(Z)=H(X,Y)$, instead of $H\big((X,Y)\big)$. We now state some basic properties of entropy for discrete random variables. These properties are well established in the literature, so we omit their proofs.
@@
In other words, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now we can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@ 2.3 Discretized entropy and properties
where for each $i$ the integer $k_{i}$ is uniquely determined by $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$. So, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in. For convenience, we often write $D_{n}$ instead of $D_{n,m}$ when the dimension is clear from the context. For any random variable $X$, the expression $D_{n}(X)$ is a discrete random variable that takes values in the set of points whose components are all dyadic rationals. Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are independent (or identically distributed, respectively).
@@
For a probability measure $\mu$, we introduce similar notation. Let $\mathcal{D}_{n}$ be the collection of all $m$-dimensional cubes $\{x\in{\mathbb{R}}^{m}:\,D_{n}(x)=k2^{-n}\}$ with $k\in\mathbb{Z}^{m}$. And define $H_{n}(\mu)=H_{n}(\mu,\mathcal{D}_{n})$. When $\mu$ is the distribution measure associated to $X$, the above two defintions coincide.
@@
Let $\mathcal{R}\subseteq{\mathbb{R}}$ and $f:\,\mathcal{R}\rightarrow{\mathbb{R}}$ be a differentiable function. Suppose that $\mathcal{R}$ is a union of $d$ intervals such that $f$ is monotonic on each of the intervals. Let $M=\sup\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$ and $m=\inf\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$. Then $f$ is weakly $(\max\{M,m^{-1}\},d)$-bi-Lipschitz on $\mathcal{R}$.
@@
When $D_{n}(f(Y)))=t$ is determined, one has $|f(Y)-t|=O_{m}(2^{-n})$ and then $|f(X)-t|\leq\kappa K+O_{m}(2^{-n})$. So $D_{n}(f(X))$ has at most $O\big((\kappa K\cdot 2^{n}+O_{m}(1))^{m}\big)$ choices. It follows that
@@
However, when $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail. In Example 7.7, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@
$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d),$
$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa)+O(\log d),$
$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa)+O(\log d).$

### v5

@@ 1.2 New discretized sum-product type estimates along dense graphs
Theorems 1.41.6 can be stated in a more analytic form, namely, by taking $A\subset[0,1]$ to be a union of closed intervals with total measure $\delta^{1-s}$. This observation has been made explicitly in [11].
For $\phi(x,y)=xy$ and a complete graph $G$, the value of $\epsilon$ has been explicitly computed and improved using a variety of techniques, see [11, 18, 19, 31] and references therein for current progress.
Note that Theorem 1.6 does not follow from the special case $B=A$, since an arbitrary subset $B$ of $A$ may not satisfy the non-concentration estimate $(B\cap I)\leq Cl^{s}(\#B)$, a phenomenon quite different from the discrete setting.
In the discrete setting, Erdős and Szemerédi [7] proposed the following strong conjecture: For every $c>0$ and $\epsilon>0$, there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, any $n$-element subset $A\subseteq\mathbb{N}$, and any graph $G$ with $n$ vertices and at least $n^{1+c}$ edges,
$(A+_{G}A)+\#(A\cdot_{G}A)\geq(\#A)^{1+c-\epsilon}.$
When $G$ is the complete graph (i.e., $c=1$), this is the famous Erdős–Szemerédi sum-product conjecture, which remains wide open. Current records can be found in [2, 32]. For arbitrary graphs $G$ with at least $n^{1+c}$ edges, Alon, Ruzsa, and Solymosi [1] refuted the strong conjecture above by constructing examples with simultaneously small sumsets and product sets along a graph. They also established that
$(A+_{G}A)+\#(A\cdot_{G}A)\gg\frac{(\#E(G))^{3/2}}{(\#A)^{7/4}}.$
Using the continuous-to-discrete adaptable-set framework developed in [14], as applications of Theorems 1.4, 1.5, and 1.6, we obtain analogous results for much more general families of polynomials in the setting of adaptable sets. Roughly speaking, if $A\subseteq[0,1]$ is a non-concentrated/adaptable $n$-point set and $G\subseteq A\times A$ is almost complete in the sense that every vertex has degree at least $n^{1-o(1)}$, then the restricted sumset $A+_{G}A$ and the restricted polynomial image $\phi_{G}(A,A)$ cannot both be small. It would be interesting to determine whether these results (Theorems 7.3, 7.4, and 7.5) can be proved independently using purely combinatorial methods.
There is a substantial parallel literature on sum-product and polynomial expansion over finite fields, especially over prime fields $\mathbb{F}_{p}$, where nontrivial lower bounds typically require the size of $A$ to be small relative to $p$. Bourgain, Katz, and Tao [3] established the first nontrivial sum-product estimate in $\mathbb{F}_{p}$, showing that $(A+A)+\#(AA)\geq(\#A)^{1+\epsilon}$ for some $\epsilon>0$ when #A=o(p)\#A=o(p). The current record lower bound is $(\#A)^{\frac{5}{4}}$ due to Mohammadi and Stevens [22].
In a more general setting, Bukh and Tsimerman [4] proved that for any polynomial $f\in\mathbb{F}_{p}[X]$ of degree $d\geq 2$ and any $A\subset\mathbb{F}_{p}$ with $A\leq\sqrt{p}$,
[…]
If $f(x)=x^{2}$, then Pham, Vinh, and de Zeeuw [29] improved the lower bound further to $(\#A)^{\frac{6}{5}}$. For a non-degenerate quadratic polynomial $\phi(x,y)$, Koh, Mojarrad, Pham, and Valculescu [16] proved that $(A+A)+\#(\phi(A,A))\gg(\#A)^{\frac{6}{5}}$. Bukh and Tsimerman [4] also established a one-sided sum-product estimate along graphs: for sets $A,B\subset\mathbb{F}_{p}$ and a graph $G\subset A\times B$ with $A,\#B\leq p^{\frac{1+\frac{1}{29}}{2}}$ and $G\geq\frac{1}{K}\#A\cdot\#B$, one has
$(A+A)+\#(A\cdot_{G}B)\gg\frac{(\#B)^{\frac{1}{29}}}{K}\#A.$
1.3 Main ideas and novelties
In this subsection, we provide an overview of the approaches for Theorems 1.2 and 1.3. The essential starting point is to find some random variable that not only links with both $X+Y$ and $\phi(X,Y)$, but also produce entropy gain.
In the case $\phi(x,y)=xy$, Máthé and O’Regan [19] use the slope (or direction) variable. Their key geometric input is a sharp lower bound for the set of directions spanned by segments joining two planar sets, due to Orponen–Shmerkin–Wang [28]. In particular, if $A\subseteq[1,2]$ has Hausdorff dimension $s$, then applying the direction-set theorem to $E=A\times A$ and $F=(-A)\times(-A)$ yields that the corresponding slope set has dimension at least $\min\{1,2s\}$. Equivalently, one obtains a large set of projective ratios
$\left\{\frac{a+b}{c+d}:a,b,c,d\in A\right\}.$
In entropy language, this gives a lower bound for the discretized entropy of the slope random variable $(X+Y)/(Z+W)$ built from i.i.d. variables $X,Y,Z,W$ supported on $A$. The second step of [19] is a Shannon-type (submodularity) inequality that controls the entropy of this slope variable in terms of the entropies of $X+Y$ and $XY$ (and the base entropy $H(X)$), and hence forces entropy growth in at least one of these simpler expressions. For a general polynomial $\phi$, adapting the methods from [19] would require direction-set estimates for a corresponding nonlinear direction map, and existing projection theorems do not provide the kind of uniform quantitative control needed in this setting. Accordingly, we work with a distance function variable instead. Our proofs contain five main steps as shown below.
(A) Entropy of distance variable (Theorem 2.38 and Theorem 5.1).
We establish a distance-entropy lower bound at dyadic scale $2^{-n}$, typically for
$Z:=|W-W^{\prime}|\qquad\text{or}\qquad Z^{2}:=|W-W^{\prime}|^{2},$
where $W^{\prime}$ is an i.i.d. copy of a planar random variable $W$ whose distribution is $(2s,C)$-Frostman and supported in a ball $B(0,R)\subseteq\mathbb{R}^{2}$.
Roughly speaking, when $s>\tfrac{1}{2}$, the distribution measure $\mu$ of $W$ has dimension $2s>1$ in the plane, i.e. it is thick enough for Fourier analytic distance-set technology to apply. Concretely, we use Mattila’s estimate (Lemma 2.36), which controls short-interval mass of the distance measure $\Delta(\mu)$ in terms of a Riesz energy:
$\Delta(\mu)([r,r+\eta])\ \ll_{t}\ r^{1/2}\,\eta^{t}\,I_{t+1/2}(\mu),$
for any $0<t\leq 1$ and $0\leq\eta\leq r$. This directly implies that the distance measure is $t$-Frostman (Lemma 2.37) with $t<2s-\frac{1}{2}$. To get an entropy gain, we would need $s<t$, which explains the required threshold of $s>\frac{1}{2}$ by Mattila’s methods.
In the range $s\leq\tfrac{1}{2}$, we instead use a distance-entropy input based on the work of Shmerkin and Wang [35]. This input is available only under a product structure: it treats distance measures between $\mu_{1}\times\nu$ and $\mu_{2}\times\nu$, and in random-variable language, it corresponds to $W=(X,Y)$ with $X$ and $Y$ independent and each marginal $(s,C)$-Frostman. The way this enters our entropy scheme is not a routine black-box reduction. Rather, we adapt the robust projection / thin-tube framework of [35] and carry out an additional conversion step that turns their robustness conclusions into a fixed-scale dyadic entropy lower bound at scale $2^{-n}$, with all losses made explicit. A central issue is quantitative bookkeeping under conditioning. In later steps, we condition on discretized events, producing families of conditional measures whose Frostman constants are highly non-uniform and may grow with the scale, in the sense that $\log C$ can be comparable to $n$. For this reason we need error terms of the form $O(\log C)$ rather than $O_{C}(1)$, so that these losses can be absorbed into entropy terms when we average over the conditioning parameters.
(B) Diagonal quadratic $\alpha X^{2}+\beta Y^{2}$ with independent random variables (Lemma 4.1).
Write $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y)$ and let $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$, with $W_{i}=(\sqrt{|\alpha|}X_{i},\sqrt{|\beta|}Y_{i})$. Then
$|W_{1}-W_{2}|^{2}=\alpha(X_{1}-X_{2})^{2}\pm\beta(Y_{1}-Y_{2})^{2}.$
So $|W_{1}-W_{2}|^{2}$ is a Lipschitz function of $(X_{1}-X_{2},\;Y_{1}-Y_{2})$ on the relevant bounded ranges. Moreover, by a polarization identity, $|W_{1}-W_{2}|^{2}$ is also a Lipschitz function of the triple
$\bigl(X_{1}+X_{2},\;Y_{1}+Y_{2},\;\alpha(X_{1}^{2}+X_{2}^{2})\pm\beta(Y_{1}^{2}+Y_{2}^{2})\bigr).$
Applying discretized submodularity (with Lipschitz error terms) yields an upper bound for $H_{n}(|W_{1}-W_{2}|^{2})$ in terms of the entropies of these sum/difference expressions. Plünnecke–Ruzsa type entropy inequalities for independent variables then control those entropies by a linear combination of $H_{n}(X+Y)$, $H_{n}(\alpha X^{2}+\beta Y^{2})$, and the marginal entropies $H_{n}(X)$ and $H_{n}(Y)$, up to logarithmic losses. Combining this with the distance-entropy lower bound for $H_{n}(|W_{1}-W_{2}|^{2})$ forces a gain in $H_{n}(X+Y)+H_{n}(\alpha X^{2}+\beta Y^{2})$.
(C) Diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)$ with independent random variables (Lemma 4.8).
A Plünnecke–Ruzsa type entropy inequality (Lemma 2.29), combined with a local weak bi-Lipschitz decomposition of $\rho_{1},\rho_{2}$ on suitable regular regions, permits a two-point differencing scheme: after conditioning on dyadic differences, one reduces degree $d$ to degree d1d-1.
More concretely, on an event where $\rho_{1}$ and $\rho_{2}$ are weakly bi-Lipschitz, Lemma 2.29 relates $H_{n}(\rho_{1}(X)+\rho_{2}(Y))$ to the two-point expression $H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))$, where $(X_{1},Y_{1}),(X_{2},Y_{2})$ are i.i.d. copies. Conditioning on $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, and using Lipschitz control, one identifies (up to logarithmic losses) this conditional entropy with
$(\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2}),\qquad(\Delta_{z}\rho)(x):=\rho(x+z)-\rho(x),$
where the difference polynomials have degree d1d-1. For gaps bounded away from 0 one applies the inductive hypothesis (with the quadratic base case given by Step (B)) to the conditional law $(X_{2},Y_{2}\,|\,D_{n}(X_{1}-X_{2})=z_{1},\ D_{n}(Y_{1}-Y_{2})=z_{2})$, which remains independent and marginally $(s_{1},\widetilde{C}_{1};\,s_{2},\widetilde{C}_{2})$-Frostman, with
$\widetilde{C}_{1}=O\!\left(\frac{C_{1}^{2}\,2^{-ns_{1}}}{P_{E}(D_{n}(X_{1}-X_{2})=z_{1})}\right),\qquad\widetilde{C}_{2}=O\!\left(\frac{C_{2}^{2}\,2^{-ns_{2}}}{P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})}\right).$
Finally, a weighted average over $(z_{1},z_{2})$ converts the $\log\widetilde{C}_{i}$ terms into entropies of the dyadic difference variables (e.g. $H_{n}(X_{1}-X_{2})$ and $H_{n}(Y_{1}-Y_{2})$), allowing one to close the induction after discarding very small differences.
There are three technical challenges in this step. First, one needs a local weak bi-Lipschitz decomposition for $\rho$ away from its critical set(s), so that entropy can be transferred between $X$ and $\rho(X)$ with only logarithmic losses in degree and coefficient size. Second, the conditional Frostman constants $\widetilde{C}_{1},\widetilde{C}_{2}$ depend on $(z_{1},z_{2})$ and can be very large; it is therefore essential that all losses enter only through $\log\widetilde{C}_{i}$, so that after averaging they become controllable entropy terms. Third, the complexity of $\Delta_{z}\rho$ deteriorates as $|z|\to 0$, so one discards very small differences in a controlled way while retaining enough mass/entropy to complete the degree reduction (see Section 4.2).
(D) Diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)$ with dependent random variables (Theorem 1.2).
A discretized entropy Balog–Szemerédi–Gowers mechanism (Theorem 3.1), allowing weakly bi-Lipschitz transformations, reduces the dependent case to a controlled form of conditional independence at dyadic scale. The losses are quantified in terms of the discretized mutual information $H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)$ and the logarithmic errors coming from conditioning on $D_{n}(\cdot)$. In particular, this dependence penalty is exactly what forces the baseline term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ to appear in Theorem 1.2:: in the jointly Frostman setting one cannot treat it as $O(1)$ and must keep it inside the main maximum.
More concretely, from $(X,Y)$ we construct auxiliary variables $(X^{\prime\prime},X^{\prime},Y^{\prime},Y^{\prime\prime})$ such that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))$. Conditioning on $(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))=(x,y)$ then yields a family of independent conditional laws $(X^{\prime\prime},Y^{\prime\prime}\,|\,(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))=(x,y))$ whose marginals are Frostman with constants
$C_{1}^{\prime}=O\!\left(\frac{C\,2^{-ns_{2}}}{P_{E}\!\big(D_{n}(Y^{\prime})=y\big)}\right),\qquad C_{2}^{\prime}=O\!\left(\frac{C\,2^{-ns_{1}}}{P_{E}\!\big(D_{n}(X^{\prime})=x\big)}\right).$
as obtained from the joint Frostman hypothesis on $(X,Y)$. Applying the independent diagonal-polynomial estimate to each conditional law and then averaging over $(x,y)$ converts the $\log C_{i}^{\prime}$ losses into entropies of the dyadic conditioning data, allowing one to close the argument.
There are two technical challenges in this step. First, the Frostman constants $C_{1}^{\prime},C_{2}^{\prime}$ vary with $(x,y)$ and may be large, so one must keep all dependence only through logarithms and then turn these into controllable entropy terms after averaging. Second, the conditioning is only on dyadic discretizations $D_{n}(\cdot)$ rather than on the underlying variables, so the usual Markov-chain interpretation in BSG theorem is not immediate: to prove Theorem 3.1 one must build an explicit coupling and verify that the required conditional i.i.d./independence relations persist after discretization, with the correct logarithmic losses.
(E) General polynomial $\phi(x,y)$ (Theorems 1.3).
The passage from diagonal forms $\rho_{1}(x)+\rho_{2}(y)$ to general forms $\phi(x,y)$ obtained by an invertible rational linear change of variables relies on controlling how discretized entropy behaves under rational linear combinations when the variables may be dependent. In particular, under conditional Frostman hypotheses, an invertible rational change of coordinates produces a pair of linear forms whose joint law remains non-concentrated (jointly Frostman with controlled constants), allowing us to reduce the general form to the diagonal core.
A key input here is a collection of discretized Ruzsa-type entropy inequalities that remain valid without independence, culminating in the quantitative comparison theorem for rational coefficients (Theorem 2.23). Concretely, for $k,\ell\in\mathbb{Q}$, Theorem 2.23 bounds the “excess entropy”
$H_{n}(kX+\ell Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
by a linear combination (with coefficients depending on $k,\ell$) of the corresponding excess for $X+Y$ and a dependence term measured by the discretized mutual information
$I_{n}(X;Y):=H_{n}(X)+H_{n}(Y)-H_{n}(X,Y),$
up to $O_{k,\ell}(1)$ errors. In particular, this dependence bookkeeping clarifies the different baseline terms in our main results: in the merely jointly Frostman regime we must work relative to $\tfrac{1}{2}(H_{n}(X)+H_{n}(Y))$, whereas in the conditionally Frostman regime it is natural to use a baseline involving $\tfrac{1}{2}H_{n}(X,Y)$, since $\tfrac{1}{2}(H_{n}(X)+H_{n}(Y))-\tfrac{1}{2}H_{n}(X,Y)=\tfrac{1}{2}I_{n}(X;Y)$.
These comparison inequalities are stable for rational coefficient relations but can fail for $\mathbb{Q}$-linearly independent real coefficients (Example 8.7). For this reason, our general-form results are stated for $\Xi\in\mathrm{GL}_{2}(\mathbb{Q})$, and the reduction to the diagonal core is carried out using rational linear algebra rather than arbitrary real changes of variables.
Finally, in the dependent setting the joint entropy $H_{n}(X,Y)$ arises naturally and must be eliminated from the final max-form lower bounds. A key new ingredient in our proof is a quantitative entropy-comparability statement (Proposition 4.10), showing that under conditional $(s,C;s,C)$-Frostman assumptions,
$H_{n}(X,Y)=H_{n}\big(X+Y,\ \phi(X,Y)\big)+O_{\phi,s,C,c}(1),$
[…]
@@ 2.1 Notations
In this paper, we denote by #S\#S the cardinality of a finite set $S$. By $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$. All logarithms we use here have base $2$. Moreover, we denote $\log_{+}x:=\max\{\log x,\,0\}$ for any $x>0$.
@@ 2.2 Entropy and basic properties
In this definition, we use the convention $0\log 0=0$. We also interpret $\mu(B)H(\mu_{B},\mathcal{A})$ as 0 when $\mu(B)=0$, even though $\mu_{B}$ is not defined in this situation. Before stating the first lemma, we introduce the following notation:
@@
Also notice that the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$. It is increasing on $0,e^{-1}]$, and decreasing on $e^{-1},1]$. With this in tow, we record some properties of entropy for partitions, whose proofs are simple exercises.
@@
(vi) For any Borel set $E$, we have $\mu(E)H(\mu_{E},\mathcal{A})\leq H(\mu,\mathcal{A}).$
@@
Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. For the random variable $Z=(X,Y)$, we denote $H(Z)=H(X,Y)$, instead of $H\big((X,Y)\big)$. We now state some basic properties of entropy for discrete random variables. These properties are well established in the literature, so we omit their proofs.
@@
In other words, when the other k1k-1 variables are given, the smaller $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now we can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@ 2.3 Discretized entropy and properties
where for each $i$ the integer $k_{i}$ is uniquely determined by $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$. That is to say, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in. For convenience, we often write $D_{n}$ instead of $D_{n,m}$ when the dimension is clear from the context. For any random variable $X$, the expression $D_{n}(X)$ is a discrete random variable that takes values in the set of points whose components are all dyadic rationals. Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are independent (or identically distributed, respectively).
@@
For a probability measure $\mu$, we introduce similar notation. Let $\mathcal{D}_{n}$ be the collection of all $m$-dimensional cubes $\{x\in{\mathbb{R}}^{m}:\,D_{n}(x)=k2^{-n}\}$ with $k\in\mathbb{Z}^{m}$. And define $H_{n}(\mu)=H_{n}(\mu,\mathcal{D}_{n})$. When $\mu$ is the distribution measure associated to $X$, the above two definitions coincide.
@@
Let $\mathcal{R}\subseteq{\mathbb{R}}$ and $f$ be a differentiable function with $domain(f)\supseteq\mathcal{R}$. Suppose that $\mathcal{R}$ is a union of $d$ intervals such that $f$ is monotonic on each of the intervals. Let $M=\sup\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$ and $m=\inf\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$. Then $f$ is weakly $(\max\{M,m^{-1}\},d)$-bi-Lipschitz on $\mathcal{R}$.
@@
When $D_{n}(f(Y))=t$ is determined, one has $|f(Y)-t|=O_{m}(2^{-n})$ and then $|f(X)-t|\leq\kappa K+O_{m}(2^{-n})$. So $D_{n}(f(X))$ has at most $O\big((\kappa K\cdot 2^{n}+O_{m}(1))^{m}\big)$ choices. It follows that
@@
However, when two real numbers $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail. In Example 8.7, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@
The last lemma in this subsection is a Plünnecke–Ruzsa type inequality with functions.
@@
$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa+\log d),$
$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa+\log d),$
$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa+\log d).$